# Motion-aware Latent Diffusion Models for Video Frame Interpolation

### v2

@@ 1 Introduction
Video frame interpolation (VFI) aims to generate intermediate frames between two consecutive video frames. It is commonly used to synthetically increase frame rate, e.g., to produce jitter-free slow-motion content [32]. VFI has also been employed in video compression [74], view synthesis [20], medical imaging [36] and animation production [61].
@@
Existing VFI methods [32, 76, 42, 52, 34] are mostly based on deep neural networks. The early deep learning methods always rely on 3D convolution [31] or RNNs [9] to model the contextual correlations of conditional neighboring frames. Inspired by the success of generative adversarial networks (GANs) in image synthesis, more recent attempts have been made to develop video frame interpolation methods by incorporating the adversarial loss. Owing to the remarkable generative capabilities of GANs, these methods exhibit significant efficacy in predicting interpolated video frames that possess natural and consistent content, holding the state-of-the-art in the literature. However, these deep learning-based methods for VFI always tend to generate unrealistic texture, artifacts and low-perceptual results. The reason is that the primary contributor to the optimization objective — and consequently the final performance of the model — remains the L1/L2-based distortion loss between their outputs and the ground-truth interpolated frames. This type of loss may not accurately reflect the perceptual quality of the interpolated videos, as described in [16]. As a result, it has been reported [15] that existing methods, while achieving high PSNR values, tend to under-perform perceptually, especially under challenging scenarios involving dynamic textures with complex motion [16].
Recently, de-noising diffusion probabilistic models (DDPMs) have garnered increasing attention and emerged as a new state-of-the-art in several areas [77, 78, 44, 25, 27]. These diffusion models have recently exhibited remarkable performance in generating realistic and perceptually-optimized images and videos, reportedly surpassing other generative models, including GANs and VAEs [17, 24]. [72, 16] are among the early methods that have explored the application of diffusion models for VFI tasks. Specifically, they address the VFI tasks as a form of conditional image generation by taking neighboring frames into the de-noising network for the target interpolated frame generation. However, these methods fail to explicitly model the inter-frame motions between the interpolated frames and the given neighboring conditional frames, a crucial factor in preventing the generation of blurred interpolated frames due to motion ambiguity. This is particularly important in complex dynamic scenes that involve intricate motions, occlusions, or abrupt changes in brightness.
@@
Extensive experiments on various VFI benchmark datasets, encompassing both low and high resolution content (up to 4K), demonstrate that our MADiff achieves the state-of-the-art performance significantly outperforming existing approaches, especially under challenging scenarios involving dynamic textures with complex motion.
Our contributions are summarized as follows:
@@
  • We demonstrate, through quantitative and qualitative experiments, that the proposed method achieves the state-of-the-art performance significantly outperforming existing approaches.
@@ 2.1 Image-to-Event Generation
Event cameras are a novel bio-inspired asynchronous sensor [21, 45] with advantages such as high dynamic range, high temporal resolution, and low power consumption [21, 4, 45]. Event cameras have the potential to provide solutions for a wide range of visually challenging scenarios and has already found extensive applications [26, 2, 58, 68, 66, 50, 5, 83, 44, 75, 43, 38]. However, acquiring event stream data is expensive and requires specific devices. Recently, several studies have exploited the event simulation from the standard camera, i.e. simulating the event stream from continuous images or video sequences. Kaiser et al. [33] generate events by simply applying a threshold on the image difference. A positive or negative event is generated depending on the pixel’s intensity difference. Pix2NVS [3] computes per-pixel luminance from conventional video frames. The technique generates synthetic events with inaccurate timestamps clustered to frame timestamps. [85, 82] propose a GAN-based method to generate the event volume by simply taking two continuous image frames as input. By modeling the correlation between continuous frames and event volumes, image-to-event generation models can capture inter-frame motion hints, which can serve as auxiliary guidance for the VFI task. In our MADiff, we directly utilized the pre-trained EventGAN [85] to extract motion hints for guiding the interpolation process.
@@ 2.2 Video Frame Interpolation
Existing VFI approaches are mostly based on deep learning, and can be generally categorized as flow-based or kernel-based. Flow-based methods rely on optical flow estimation to generate interpolated frames. To obtain the optical flow from input frames to the non-existent middle frame (or the other way around), some methods [32, 51, 52, 60] assume certain motion types to infer the intermediate optical flows using the flows between two input frames, while others [46, 76, 54, 55, 48, 41] directly estimate the intermediate flows. On the other hand, kernel-based methods argue that optical flows can be unreliable in dynamic texture scenes, so they predict locally adaptive convolution kernels to synthesize output pixels, allowing more flexible many-to-one mapping. While earlier methods [53] in this class predict fixed-size kernels, more recent ones [42, 18, 7, 8, 12] tend to adopt deformable convolution [11] kernels. Other than these two categories, there are also attempts that combine flow-based and kernel-based methods to perform end-to-end frame synthesis [10, 34].
@@ 5.2 Experimental Setup
Following [16], we adopt a perceptual image quality metric LPIPS [81] and a VFI quality metric FloLPIPS [13] for evaluation. These metrics have shown superior correlation with human judgments of VFI quality compared to commonly used quality measurements, PSNR and SSIM [73]. We also evaluate FID [22] which measures the similarity between the distributions of target and ground-truth frames. FID was previously used as a perceptual metric for video compression [80], enhancement [79] and colorization [35]. And we also provide benchmark results based on PSNR and SSIM in Appendix B, noting that these are limited in reflecting the perceptual quality of interpolated content [15] and are not the focus of this paper.

### v3

@@ 1 Introduction
Video frame interpolation (VFI) aims to generate intermediate frames between two consecutive video frames. It is commonly used to increase the frame resolution, e.g., to produce slow-motion content [30]. VFI has also been applied to video compression [76], video generation [21] and animation production [62].
@@
Existing VFI methods [30, 78, 39, 50, 32] are mostly based on deep neural networks. The early deep learning methods always rely on 3D convolution [29] or RNNs [9] to model the contextual correlations of conditional neighboring frames. Inspired by the success of generative adversarial networks (GANs) in image synthesis [59, 55, 52, 22], more recent attempts have been made to develop video frame interpolation methods by incorporating the adversarial loss. Owing to the remarkable generative capabilities of GANs, these methods exhibit significant efficacy in predicting interpolated video frames that possess natural and consistent content, holding the state-of-the-art in the literature. However, these deep learning-based methods for VFI always tend to generate unrealistic texture, artifacts and low-perceptual results. The reason is that the primary contributor to the optimization objective — and consequently the final performance of the model — remains the L1/L2-based distortion loss between their outputs and the ground-truth interpolated frames [15]. This type of loss may not accurately reflect the perceptual quality of the interpolated videos. Consequently, [14] indicates that current techniques, despite attaining elevated PSNR scores, often fall short in terms of perceptual quality, particularly in demanding situations characterized by dynamic textures and complex motion patterns.
Recently, de-noising diffusion probabilistic models (DDPMs) have been gaining widespread interest and have risen to prominence as state-of-the-art solutions across various areas [66, 20, 21, 79, 23, 25]. These diffusion models have demonstrated exceptional capabilities in creating realistic and perceptually-enhanced images and videos, reportedly outperforming other generative models. [74, 15] are among the early methods that have explored the application of diffusion models for VFI tasks. Specifically, they address the VFI tasks as a form of conditional image generation by taking neighboring frames into the de-noising network for the target interpolated frame generation.
However, these methods fail to explicitly model the inter-frame motions between the interpolated frames and the given neighboring conditional frames, a crucial factor in preventing the generation of blurred interpolated frames due to motion ambiguity. This is particularly important in complex dynamic scenes that involve intricate motions, occlusions, or abrupt changes in brightness.
@@
Extensive experiments on various VFI benchmark datasets, encompassing both low and high resolution content (up to 4K), demonstrate that our MADiff achieves the state-of-the-art performance significantly outperforming existing approaches, especially under challenging scenarios involving dynamic textures with complex motion. Our contributions are summarized as follows:
@@
  • We demonstrate, through quantitative and qualitative experiments, that the proposed method achieves the state-of-the-art performance outperforming existing approaches.
@@ 2.1 Image-to-Event Generation
Event cameras are a novel bio-inspired asynchronous sensor [18, 43] with advantages such as high dynamic range, high temporal resolution, and low power consumption [18, 4, 43]. Event cameras have the potential to provide solutions for a wide range of visually challenging scenarios and has already found extensive applications [2, 58, 70, 68, 48, 5, 84, 42, 41, 77, 40, 35, 24]. However, acquiring event stream data is expensive and requires specific devices. Recently, several studies have exploited the event simulation from the standard camera, i.e. simulating the event stream from continuous images or video sequences. Kaiser et al. [31] generate events by simply applying a threshold on the image difference. A positive or negative event is generated depending on the pixel’s intensity difference. Pix2NVS [3] computes per-pixel luminance from conventional video frames. The technique generates synthetic events with inaccurate timestamps clustered to frame timestamps. [86, 83] propose a GAN-based method to generate the event volume by simply taking two continuous image frames as input. By modeling the correlation between continuous frames and event volumes, image-to-event generation models can capture inter-frame motion hints, which can serve as auxiliary guidance for the VFI task. In our MADiff, we directly utilized the pre-trained EventGAN [86] to extract motion hints for guiding the interpolation process.
@@ 2.2 Video Frame Interpolation
Current approaches to video frame interpolation (VFI) that leverage deep learning can typically be divided into two main categories: those that are based on flow estimation and those that rely on kernel methods. Existing flow-based methods [30, 49, 50, 61, 44, 78, 53, 54, 46, 38] primarily generate intermediate frames by analyzing the pixel motion between consecutive frames. Specifically, these methods use optical flow estimation to capture the movement of objects within the scene, and then synthesize intermediate frames based on these estimations. These methods can be founded on various assumptions and algorithms, such as assuming uniform motion within the scene or employing machine learning models to predict the optical flow field, thereby enhancing the accuracy and visual quality of the interpolated frames. On the other hand, Existing kernel-based video frame interpolation methods [51, 39, 16, 7, 8, 11] typically utilize kernel functions to estimate the relationships between pixels across different frames, generating intermediate frames by performing a weighted average of pixel values based on known frames. These methods can flexibly adapt to various motion patterns in the scene by selecting appropriate kernel sizes and shapes, thus effectively smoothing the visual transitions between frames while preserving details. In addition to these two main categories, there are also some methods integrate flow-based and kernel-based approaches to better synthesis performance [10, 32].
@@ 5.2 Experimental Setup
Following [15], we adopt a perceptual image quality metric LPIPS [82], FloLPIPS [12] for performance evaluation. These metrics exhibit a stronger correlation with human assessments of VFI quality in contrast to the traditionally utilized quality metrics, such as PSNR and SSIM [75]. We also evaluate FID [19] which measures the similarity between the distributions of interpolated and ground-truth frames. FID was previously used as a perceptual metric for video compression [81], enhancement [80] and colorization [33]. We also provide benchmark results based on PSNR and SSIM in the Appendix, noting that these are limited in reflecting the perceptual quality of interpolated content [14] and are therefore not the focus of this paper.